[Brahms-l] Shift report 20040330 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Tue Mar 30 2004 - 07:51:55 EST
------------------------------------------------------------
Date: Tue Mar 30 07:51:54 2004, Shift: 00:00-08:00
Supervisor: BB, with: 
Log book#15 pages 102 - 102

Shift Summary:
------------------------------------------------------------
1. Beam for 3.25 hours. Two unplanned dumps.

2.All runs with MRS at 40 degrees 6kg B
                FS at 4 degrees 1/4 B
 Run     trigger 3  trigger 6
11434       93k       137k
11437       68k       100k
11438       37k        58k
11439       20k        29k

3. No Detector problems. However, experienced 
   daq hangups at start of runs. Restarting 
   RunServer and IoServer worked.(See FAQ)
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue Mar 30 07:52:40 2004

This archive was generated by hypermail 2.1.8 : Tue Mar 30 2004 - 07:53:00 EST